OpenStudy (anonymous):

8x^-3y^-1 over -2x^6y^-4

OpenStudy (anonymous):

its -(4 x^3)/y^5 or (-4 x^3)/y^5

OpenStudy (johnweldon1993):

\[\large \frac{8x^{-3}y^{-1}}{-2x^6y^{-4}}\]

OpenStudy (anonymous):

−4x^−3y−6y4− that is what I got.?

OpenStudy (johnweldon1993):

If we use the rule for dividing exponents with the same base...we just subtract the exponents...so we have in terms of the 'x' term....we have\[\large \frac{x^{-3}}{x^6} = x^{-3 - 6} = x^{-9}\] and now interms of the 'y' we have \[\large \frac{y^{-1}}{y^{-4}} = y^{-1 - (-4)} = y^{-1 + 4} = y^{3}\] so we have \[\large -4x^{-9}y^{3}\] or written as \[\large \frac{-4y^3}{x^9}\]

OpenStudy (anonymous):

ill try it

OpenStudy (anonymous):

I WAS RIGHT!!!!!!! WOHOOO LOL anyhow thx john